# Lifelong Learning with Behavior Consolidation for Vehicle Routing

### v3

@@ 1 Introduction
Existing studies on lifelong learning for neural VRP solvers (Li et al., 2025; Feng et al., 2025) are largely confined to highly specific scenario settings: tasks solely differing in scale or distance metric, the task order is fixed and known, and the generation of new problem instances is controllable. Furthermore, they are mainly dedicated to the solver’s performance at the end of the lifelong learning process after being trained on all training tasks. As lifelong learning is an ongoing process (possibly without any end), the solver’s performance at each time point of the lifelong learning process is also important. Their proposed methods rely on actively generating and training on new problem instances from previous tasks, making them inapplicable when the generation of new instances is uncontrollable, as is the case studied in this work.
@@ 2.2 Cross-distribution Cross-scale Generalization
However, these works mainly focus on one-off training on a predefined task set. New tasks are addressed with zero-shot generalization or fine-tuning the trained solver on each of them independently. The potential of sequential knowledge accumulation from multiple new tasks is overlooked, together with the impact of catastrophic forgetting. As a result, their performance in learning to solve sequentially arising tasks is still limited. In practice, tasks arrive continuously, and past tasks can inform future learning. This motivates the need for lifelong neural VRP solvers that can incrementally learn from a sequence of diverse tasks.
@@ 2.3 Lifelong Learning
To the best of our knowledge, only two existing works have studied lifelong learning for neural VRP solver (Li et al., 2025; Feng et al., 2025). However, their studies focused on rather restricted scenarios: (i) tasks differ only in scale (Li et al., 2025; Feng et al., 2025) or distance metric (Feng et al., 2025), (ii) task order is fixed and known under strong assumptions (e.g., task scale gradually increases), and (iii) problem instance generation is controllable so that new instances of previously learned tasks can always be actively generated, making these methods inapplicable to scenarios with uncontrollable instance streams. Studies of more generic and practical scenarios are desired.
@@ 3 Lifelong Learning Router with Behavior Consolidation
Figure 2: LLR-BC in the lifelong learning scenario where new tasks with different distributions and scales sequentially arise over time. $T_{t}$: the task at time $t$. $\pi_{\theta}$: solver with model parameters $\theta$. $\mathcal{B}$: experience buffer. $\mathcal{J}$: generated problem instances. $\{\tau\}$: problem solving trajectories. $\mathcal{E}$: experiences sampled from the buffer. $w_{\mathcal{E}}$: weights of experiences. $a$: action from action space $\mathcal{A}$, i.e., node to visit. $s$: state. var: variance. $\mathcal{P}$: probability distribution over actions.
@@ 3.1 Experience Replay for Routing
While existing methods (Li et al., 2025; Feng et al., 2025) take one whole problem instance of previous tasks as an experience, we consider experiences at finer granularity. We follow the commonly studied Markov decision process (MDP) formulation of constructive neural VRP solver (Kwon et al., 2020; Zhou et al., 2023; Bogyrbayeva et al., 2024), where each state corresponds to a partial solution composed of visited nodes, and an action is to select one node to visit next. In LLR-BC, each experience $e=\langle s,\mathcal{P}\rangle$ consists of a state $s$ and the solver’s behavior $\mathcal{P}$ on the state $s$. A state $s$ represents the current partial solution for solving an instance. The behavior $\mathcal{P}$ is the probability distribution of selecting each node as the next visit. The detailed solver design, including the handling of infeasible actions and the encoding of problems and solutions, follows any underlying base neural solver used. We use the probability distribution over all nodes rather than the single selected node to represent the solver’s behavior, as it captures richer information about the learned routing strategy (Rusu et al., 2016; Buzzega et al., 2020).
To consolidate high-quality behaviors, LLR-BC buffers experiences only during the final epoch of each task, as in the final epoch the solver is well trained on the task and can provide high-quality behaviors. At each model update step, a random set of experiences $\mathcal{E}$ is sampled from buffer $\mathcal{B}$ for behavior consolidation. A detailed discussion of the additional memory usage from the buffer can be found in Appendix A.5.
@@ 3.3 Decision-seeking Behavior Consolidation (DsBC)
Existing methods (Li et al., 2025; Feng et al., 2025) replay experiences by generating, solving, and learning instances of previous tasks, ignoring the instructive information contained in historical behaviors. In contrast, by minimizing the weighted sum of the difference between the current model’s behaviors and the buffered behavior at buffered states, LLR-BC encourages the consolidation of prior behavior patterns, thereby preserving knowledge and mitigating catastrophic forgetting.
The Kullback–Leibler divergence (KLD) $D_{KL}(P||Q)$, where $P$ and $Q$ are the teacher and leaner probability distribution respectively, is widely adopted for knowledge distillation of neural VRP solvers (Bi et al., 2022; Li et al., 2025; Zheng et al., 2025). However, recent studies (Kaplanis et al., 2019; Wang et al., 2024a; Qi et al., 2026) have shown that using reverse KLD (RKLD) $D_{RKL}(P||Q)=D_{KL}(Q||P)$ (detailed explanation is in Appendix A.3), could lead to better learning performance. While minimizing KLD pushes the learner distribution to spread probability mass across all modes of the given teacher distribution, minimizing RKLD leads to mode-seeking, where the learner tends to concentrate on the teacher’s highest-probability actions, while still maintaining relatively good alignment with the overall distribution (Chan et al., 2022; Wu et al., 2025). During problem-solving of constructive neural VRP solvers, the action (node) with the highest probability is typically chosen to extend the route. To address forgetting, that could manifest in significantly longer routes on previously seen tasks, it is crucial to preserve these highest-probability decisions. Therefore, instead of KLD, LLR-BC employs the RKLD to measure the behavioral difference in a decision-seeking way. To consolidate sampled behaviors, LLR-BC minimizes the following behavior consolidation loss term, considering the RKLD (Wang et al., 2024a; Qi et al., 2026):
@@ 4 Experiments
To answer the questions, we simulate multiple lifelong learning scenarios on CVRP and TSP, and compare LLR-BC against widely used and representative lifelong learning baselines across diverse metrics [^1] Code will be released upon acceptance of this paper.. Most experiments are conducted based on POMO (Kwon et al., 2020), given its concise design and broad applicability. To verify the applicability of LLR-BC, we further evaluate LLR-BC on two representative and state-of-the-art constructive neural solvers designed for good cross-distribution cross-scale performance, i.e., Omni (Zhou et al., 2023) and INViT (Fang et al., 2024).
@@ 4.2 Stability and Plasticity
We analyze the stability by comparing the AF and AMF values over all five task orders, as shown in Table 1. Li (inter) and Li (intra) (Li et al., 2025), perform better in mitigating forgetting than LLR-BC. This is because half of their training resources (in terms of epochs) on the new task are used to learn purely from instances of previous tasks. Although they address forgetting well, they do not learn much from new tasks and perform poorly in terms of AP and APl. LLR-BC achieves substantially lower AF and AMF values than the rest of the compared methods, demonstrating its strong stability for effectively mitigating catastrophic forgetting. Figure 3 shows the average solution distance on seen tasks during lifelong learning of task order 1. At the start of each task, most baselines exhibit a sharp performance drop on previously learned tasks. Li (inter) and Li (intra) learn new tasks slowly. Feng suffers from low stability in maintaining learned knowledge. In contrast, LLR-BC mitigates forgetting and maintains consistently lower distances with greater stability.
@@ 5 Conclusions
We propose LLR-BC, a lifelong learning framework for neural VRP solvers, considering tasks with varying problem scales and distributions that arrive sequentially and need to be learned. We introduce two core components: CaEW, which emphasizes low-confidence behaviors, and DsBC, which preserves learned behavior effectively in a decision-seeking way. Experiments across diverse orders of tasks with varying distributions and scales demonstrate that LLR-BC consistently outperforms baselines in terms of performance in solving learned tasks, resistance to catastrophic forgetting, learning efficiency on new tasks, and zero-shot generalization. Ablation and sensitivity studies confirm the effectiveness of the core designs of LLR-BC and the framework’s robustness to hyperparameter settings. LLR-BC is both model- and RL algorithm-agnostic, and can be integrated with various neural VRP solvers. Experiments on different base neural solvers verify LLR-BC’s applicability.
@@
All algorithmic details (cf. Section 3 and Appendix A), training protocols (cf. Section 4 and Appendix C), and evaluation metrics (cf. Section 4 and Appendix C.2) are described in the main paper and further elaborated in the Appendix. For empirical studies, we provide a detailed description of the datasets and instance generation pipeline (cf. Appendix B). Hyperparameters and implementation details for all baselines are also reported in Section 4 and Appendix C. Upon acceptance, we will release our code and scripts for reproducing our experiments, including instructions for running and data preparation. Together, these resources enable independent researchers to replicate our results and build upon our contributions.
@@ C.1 Baselines
To adapt Li (inter), Li (intra), and Feng to our focused scenarios where the generation of new problem instances is uncontrollable, we maintain one instance buffer for each seen task instead of generating new instances for their experience replay, so that each buffer contains the instances generated during learning the corresponding task. Buffers are updated with reservoir sampling. The buffer size of each task is 200, so that there is a total of 1000 buffered instances that can be used for learning the 6th tasks, aligning with the buffer size of LLR-BC. We set $\alpha=0.5$ arbitrarily for Li (inter), Li (intra) and Feng, as we did not find suggested value or method for setting the value of $\alpha$.. Feng contains a learnable model parameter $\mathbf{B}$ with fixed size $n\times n$ ($n$ is the problem scale). We did not find the description about the way to handle $\mathbf{B}$ under scenarios where problem scale changes, so we removed $\mathbf{B}$ to make it applicable to our scenarios. To adapt LiBOG in VRP tasks, we directly embed its two key modules, i.e., the intra-task and inter-task consolidation terms, into POMO for lifelong learning. We set weight of the consolidation term to $10$ for EWC and LiBOG, and the weight of the intra-task consolidation term to 1e41e-4 for LiBOG. For all runs, the solver learns the first task from an identical randomly generated model. As Restart learns each task independently under any task order without being affected by the order, we run Restart once on each task and calculate metrics under different task orders. Therefore, Restart’s APl owns not std. value. As in most cases, POMO can complete the learning on each single task for 200 epochs (Kwon et al., 2020), we let each method train for 200 epochs on each task. To ensure comparable training time across tasks of different problem scales, we vary the number of training problems sampled per epoch: 10,000 for tasks of scale 20, 4,000 for scale 50, and 2,000 for scale 100, with batch sizes of 64, 32, and 16, respectively. Following  Kwon et al. (2020), during training, we sample a batch of problem instances and run $N$ parallel solving processes per instance, each corresponding to a different starting node, where $N$ is the problem scale. This results in $N^{2}$ parallel trajectories per batch, yielding $N^{2}$ independent states and behaviors per experience batch. The size of the experience batch varies across tasks with different problem scales. We consider this reasonable, as larger-scale tasks are typically more complex and thus require larger experience batches to effectively retain learned knowledge. We experimented with buffering and sampling at the level of individual experiences rather than experience batches. However, when using GPU-based parallelization, operating at the batch level significantly reduces computation time without yielding significant performance differences.
@@ C.3 Ablation Study Details
For instance-based buffering, we conduct batch-level buffering, set the buffer size to 20, so that the expected buffered steps are more than the buffer size in the step-based buffer (20 * $\frac{20+20+50+50+100+100}{6}\approx 1133\textgreater 1000$). And 1 (batch of) instances will be sampled from the buffer for each consolidation, i.e., $\mathcal{E}=1$. With the similar buffer size (in terms of buffered steps (state-behavior pairs) ), instance-based design reduces the instance-diversity of the buffer, potentially leading to inferior performance. The experimental results indicate that our state-based buffer design outperforms the instance-based buffer design, supporting the its advantage.
Though reservoir sampling gives each experience an equal probability of being kept in the buffer, since the learning processes of different tasks generate different numbers of experiences, the experiences in the buffer are not uniformly added from all tasks. For tasks with a larger problem scale, more experiences from these tasks are buffered. For one run on task order 1 with buffer size 1000, after learning all tasks, the numbers of experiences in the buffer from each task are: E:162, C:255, G:141, U:78, R:77, GM:287. The ratio of the numbers of buffered experiences is highly correlated with the ratio of the problem scales. We suggest this is reasonable, as tasks with larger scale may be more difficult and need more experiences for consolidating their knowledge and addressing catastrophic forgetting during later learning. After training all tasks with the rescaling version -Res, the numbers of experiences in the buffer from each task are: E:148, C:148, G:144, U:199, R:208, GM:153. Though the buffer is more balanced, it performs (cf. Table 3) very similarly to (or even slightly worse than) the original LLR-BC, suggesting the effectiveness of our design.
@@ C.4 Other Details
Our experiments are conducted on a GPU cluster utilizing a single Nvidia A100 GPU per run. Each lifelong learning method implemented based on POMO [^2] https://github.com/yd-kwon/POMO (Kwon et al., 2020) and Omni [^3] https://github.com/RoyalSkye/Omni-VRP (Zhou et al., 2023) builds upon the official codes released by the original authors.

### v4

@@ 1 Introduction
Existing studies on lifelong learning for neural VRP solvers (Li et al., 2024; Feng et al., 2025) are largely confined to highly specific scenario settings: tasks solely differing in scale or distance metric, the task order is fixed and known, and the generation of new problem instances is controllable. Furthermore, they are mainly dedicated to the solver’s performance at the end of the lifelong learning process after being trained on all training tasks. As lifelong learning is an ongoing process (possibly without any end), the solver’s performance at each time point of the lifelong learning process is also important. Their proposed methods rely on actively generating and training on new problem instances from previous tasks, making them inapplicable when the generation of new instances is uncontrollable, as is the case studied in this work.
@@ 2.2 Cross-distribution Cross-scale Generalization
However, existing studies largely rely on one-off training on a predefined task set. New tasks are typically handled via zero-shot generalization or independent fine-tuning, without leveraging knowledge accumulated across new tasks, and with limited consideration of catastrophic forgetting. Consequently, their effectiveness in sequential task learning remains limited. In realistic settings, tasks arrive continuously and previously learned tasks can inform future adaptation, motivating the development of lifelong neural VRP solvers that incrementally learn from a stream of diverse tasks.
@@ 2.3 Lifelong Learning
To the best of our knowledge, only two existing works have studied lifelong learning for neural VRP solver (Li et al., 2024; Feng et al., 2025). However, their studies focused on rather restricted scenarios: (i) tasks differ only in scale (Li et al., 2024; Feng et al., 2025) or distance metric (Feng et al., 2025), (ii) task order is fixed and known under strong assumptions (e.g., task scale gradually increases), and (iii) problem instance generation is controllable so that new instances of previously learned tasks can always be actively generated, making these methods inapplicable to scenarios with uncontrollable instance streams. Studies of more generic and practical scenarios are desired.
@@ 3 Lifelong Learning Router with Behavior Consolidation
Figure 2: LLR-BC in the lifelong learning scenario where new tasks with different distributions and scales sequentially arise over time. $T_{t}$: the task at time $t$. $\pi_{\theta}$: solver with model parameters $\theta$. $\mathcal{B}$: experience buffer. $\mathcal{J}$: problem instances. $\{\tau\}$: problem solving trajectories. $\mathcal{E}$: experiences. $w_{\mathcal{E}}$: weights of experiences. $a$: action, i.e., node to visit. $s$: state. var: variance. $\mathcal{P}$: behavior.
@@ 3.1 Experience Replay for Routing
While existing methods (Li et al., 2024; Feng et al., 2025) take one whole problem instance of previous tasks as an experience, we consider experiences at finer granularity. We follow the commonly studied Markov decision process (MDP) formulation of constructive neural VRP solver (Kwon et al., 2020; Zhou et al., 2023; Bogyrbayeva et al., 2024), where each state corresponds to a partial solution composed of visited nodes, and an action is to select one node to visit next. In LLR-BC, each experience $e=\langle s,\mathcal{P}\rangle$ consists of a state $s$ and the solver’s behavior $\mathcal{P}$ on the state $s$. A state $s$ represents the current partial solution for solving an instance. The behavior $\mathcal{P}$ is the probability distribution of selecting each node as the next visit. The detailed solver design, e.g., the encoding of problems and solutions, follows the underlying base neural solver used. We use the probability distribution over all nodes rather than the single selected node to represent the solver’s behavior, as it captures richer information about the learned routing strategy (Rusu et al., 2016; Buzzega et al., 2020).
To consolidate high-quality behaviors, LLR-BC buffers experiences only during the final epoch of each task, when the solver is expected to be well-trained on the task. At each model update step, a random set of experiences $\mathcal{E}$ is sampled from buffer $\mathcal{B}$ for behavior consolidation. A detailed discussion of the additional memory usage from the buffer can be found in Appendix A.5.
@@ 3.3 Decision-seeking Behavior Consolidation (DsBC)
Existing methods (Li et al., 2024; Feng et al., 2025) replay experiences by generating, solving, and learning instances of previous tasks, ignoring the instructive information contained in historical behaviors. In contrast, by minimizing the weighted sum of the difference between the current model’s behaviors and the buffered behavior at buffered states, LLR-BC encourages the consolidation of prior behavior patterns, thereby preserving knowledge and mitigating catastrophic forgetting.
The Kullback–Leibler divergence (KLD) $D_{KL}(P||Q)$, where $P$ and $Q$ are the teacher and leaner probability distribution respectively, is widely adopted for knowledge distillation of neural VRP solvers (Bi et al., 2022; Li et al., 2024; Zheng et al., 2025). However, recent studies (Kaplanis et al., 2019; Wang et al., 2024a; Qi et al., 2025) have shown that using reverse KLD (RKLD) $D_{RKL}(P||Q)=D_{KL}(Q||P)$ (detailed explanation is in Appendix A.3), could lead to better learning performance. While minimizing KLD pushes the learner distribution to spread probability mass across all modes of the given teacher distribution, minimizing RKLD leads to mode-seeking, where the learner tends to concentrate on the teacher’s highest-probability actions, while still maintaining relatively good alignment with the overall distribution (Chan et al., 2022; Wu et al., 2025). During problem-solving of constructive neural VRP solvers, the action (node) with the highest probability is typically chosen to extend the route. To address forgetting, that could manifest in significantly longer routes on previously seen tasks, it is crucial to preserve these highest-probability decisions. Therefore, instead of KLD, LLR-BC employs the RKLD to measure the behavioral difference in a decision-seeking way. To consolidate sampled behaviors, LLR-BC minimizes the following behavior consolidation loss term, considering the RKLD (Wang et al., 2024a; Qi et al., 2025):
@@ 4 Experiments
To answer the questions, we simulate multiple lifelong learning scenarios on CVRP and TSP, and compare LLR-BC against widely used and representative lifelong learning baselines across diverse metrics [^1] Our implementation is available in https://github.com/PeiJY/LLR-BC.. Most experiments are conducted based on POMO (Kwon et al., 2020), given its concise design and broad applicability. To verify the applicability of LLR-BC, we further evaluate LLR-BC on two representative and state-of-the-art constructive neural solvers designed for good cross-distribution cross-scale performance, i.e., Omni (Zhou et al., 2023) and INViT (Fang et al., 2024).
@@ 4.2 Stability and Plasticity
We analyze the stability by comparing the AF and AMF values over all five task orders, as shown in Table 1. Li (inter) and Li (intra) (Li et al., 2024), perform better in mitigating forgetting than LLR-BC. This is because half of their training resources (in terms of epochs) on the new task are used to learn purely from instances of previous tasks. Although they address forgetting well, they do not learn much from new tasks and perform poorly in terms of AP and APl. LLR-BC achieves substantially lower AF and AMF values than the rest of the compared methods, demonstrating its strong stability for effectively mitigating catastrophic forgetting. Figure 3 shows the average solution distance on seen tasks during lifelong learning of task order 1. At the start of each task, most baselines exhibit a sharp performance drop on previously learned tasks. Li (inter) and Li (intra) learn new tasks slowly. Feng suffers from low stability in maintaining learned knowledge. In contrast, LLR-BC mitigates forgetting and maintains consistently lower distances with greater stability.
@@ 5 Conclusions
We propose LLR-BC, a lifelong learning framework for neural VRP solvers, considering learning tasks with varying problem scales and distributions that arrive sequentially. We introduce two core components: CaEW, which emphasizes low-confidence behaviors, and DsBC, which preserves learned behavior effectively in a decision-seeking way. Experiments across diverse orders of tasks with varying distributions and scales demonstrate that LLR-BC consistently outperforms baselines in terms of performance in solving learned tasks, resistance to catastrophic forgetting, learning efficiency on new tasks, and zero-shot generalization. Ablation and sensitivity studies confirm the effectiveness of the core designs of LLR-BC and the framework’s robustness to hyperparameter settings. LLR-BC is both model- and RL algorithm-agnostic, and can be integrated with various neural VRP solvers. Experiments on different base neural solvers verify LLR-BC’s applicability.
@@
Acknowledgement
This work was supported by the internal grants of the Lingnan University, Hong Kong SAR, China. It was also supported in part by the New Zealand MBIE Endeavour Smart Ideas Grant under Contract RTVU2305 and MBIE SSIF Fund on Data Science Programme under contract RTVU1914. Xin Yao was supported by an internal grant from Lingnan University, Hong Kong SAR, China
@@ Acknowledgement
All algorithmic details (cf. Section 3 and Appendix A), training protocols (cf. Section 4 and Appendix C), and evaluation metrics (cf. Section 4 and Appendix C.2) are described in the main paper and further elaborated in the Appendix. For empirical studies, we provide a detailed description of the datasets and instance generation pipeline (cf. Appendix B). Hyperparameters and implementation details for all baselines are also reported in Section 4 and Appendix C. We also release our code and scripts for reproducing our experiments, including instructions for running and data preparation. Together, these resources enable independent researchers to replicate our results and build upon our contributions.
@@ C.1 Baselines
To adapt Li (inter), Li (intra), and Feng to our focused scenarios where the generation of new problem instances is uncontrollable, we maintain one instance buffer for each seen task instead of generating new instances for their experience replay, so that each buffer contains the instances generated during learning the corresponding task. Buffers are updated with reservoir sampling. The buffer size of each task is 200, so that there is a total of 1000 buffered instances that can be used for learning the 6th tasks, aligning with the buffer size of LLR-BC. We set $\alpha=0.5$ arbitrarily for Li (inter), Li (intra) and Feng, as we did not find a suggested value or method for setting the value of $\alpha$. Feng contains a learnable model parameter $\mathbf{B}$ with fixed size $n\times n$ ($n$ is the problem scale). We did not find the description about the way to handle $\mathbf{B}$ under scenarios where problem scale changes, so we removed $\mathbf{B}$ to make it applicable to our scenarios. To adapt LiBOG in VRP tasks, we directly embed its two key modules, i.e., the intra-task and inter-task consolidation terms, into POMO for lifelong learning. We set the weight of the consolidation term to $10$ for EWC and LiBOG, and the weight of the intra-task consolidation term to 1e41e-4 for LiBOG. For all runs, the solver learns the first task from an identical randomly generated model. As Restart learns each task independently under any task order without being affected by the order, we run Restart once on each task and calculate metrics under different task orders. Therefore, Restart’s APl owns not std. value. As in most cases, POMO can complete the learning on each single task for 200 epochs (Kwon et al., 2020), we let each method train for 200 epochs on each task. To ensure comparable training time across tasks of different problem scales, we vary the number of training problems sampled per epoch: 10,000 for tasks of scale 20, 4,000 for scale 50, and 2,000 for scale 100, with batch sizes of 64, 32, and 16, respectively. Following  Kwon et al. (2020), during training, we sample a batch of problem instances and run $N$ parallel solving processes per instance, each corresponding to a different starting node, where $N$ is the problem scale. This results in $N^{2}$ parallel trajectories per batch, yielding $N^{2}$ independent states and behaviors per experience batch. The size of the experience batch varies across tasks with different problem scales. We consider this reasonable, as larger-scale tasks are typically more complex and thus require larger experience batches to effectively retain learned knowledge. We experimented with buffering and sampling at the level of individual experiences rather than experience batches. However, when using GPU-based parallelization, operating at the batch level significantly reduces computation time without yielding significant performance differences.
@@ C.3 Ablation Study Details
For instance-based buffering, we conduct batch-level buffering, set the buffer size to 20, so that the expected buffered steps are more than the buffer size in the step-based buffer (20 * $\frac{20+20+50+50+100+100}{6}\approx 1133\textgreater 1000$). And 1 (batch of) instances will be sampled from the buffer for each consolidation, i.e., $\mathcal{E}=1$. With the similar buffer size (in terms of buffered steps (state-behavior pairs) ), instance-based design reduces the instance-diversity of the buffer, potentially leading to inferior performance. The experimental results indicate that our state-based buffer design outperforms the instance-based buffer design, supporting its advantage.
Though reservoir sampling gives each experience an equal probability of being kept in the buffer, since the learning processes of different tasks generate different numbers of experiences, the experiences in the buffer are not uniformly added from all tasks. For tasks with a larger problem scale, more experiences from these tasks are buffered. For one run on task order 1 with buffer size 1000, after learning all tasks, the numbers of experiences in the buffer from each task are: E:162, C:255, G:141, U:78, R:77, GM:287. The ratio of the numbers of buffered experiences is highly correlated with the ratio of the problem scales. We suggest this is reasonable, as tasks with larger scales may be more difficult and need more experiences for consolidating their knowledge and addressing catastrophic forgetting during later learning. After training all tasks with the rescaling version -Res, the numbers of experiences in the buffer from each task are: E:148, C:148, G:144, U:199, R:208, GM:153. Though the buffer is more balanced, it performs (cf. Table 3) very similarly to (or even slightly worse than) the original LLR-BC, suggesting the effectiveness of our design.
@@ C.4 Other Details
Our experiments are conducted on a GPU cluster utilizing a single Nvidia A100 GPU per run. Each lifelong learning method implemented based on POMO [^2] https://github.com/yd-kwon/POMO (Kwon et al., 2020), Omni [^3] https://github.com/RoyalSkye/Omni-VRP (Zhou et al., 2023) , and INViT [^4] https://github.com/Kasumigaoka-Utaha/INViT (Fang et al., 2024) builds upon the official codes released by the original papers.